OpenStudy (anonymous):

Find the tangent line to the curve \[y=\cos(x+y)\] at the point \[(\Pi/2,0)\]

OpenStudy (solomonzelman):

You need to find dy/dx of the curve

OpenStudy (solomonzelman):

Do you have any trouble differentiating on both sides? (chain rule twice on the right side - once for the part inside the cosine - second chain for the y (the y') )

OpenStudy (solomonzelman):

Ok, what is the derivative of y?

OpenStudy (anonymous):

Would it be y'?

OpenStudy (solomonzelman):

yes exactly

OpenStudy (anonymous):

So it's -sin(x+y) (y')

OpenStudy (solomonzelman):

y is a function of x, (just as f(x) is) The derivative of f(x) is f'(x), AND the derivative of y is y'.

OpenStudy (solomonzelman):

the derivative of x is 1, not 0

OpenStudy (solomonzelman):

(Nice username btw, I like that)

OpenStudy (anonymous):

Oh okay, I forgot the 1. \[-\sin(x+y)\times(1+y')=f'(x)\]

OpenStudy (solomonzelman):

yes, y' = -sin(x+y) × (1+y')

OpenStudy (anonymous):

Thanks!

OpenStudy (solomonzelman):

Ok, now you need to isolate the y'

OpenStudy (solomonzelman):

(y' is same as dy/dx)

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle y' = -\sin(x+y) \times (1+y') }\) \(\large\color{#000000 }{ \displaystyle y' = -\sin(x+y) -y'\sin(x+y) }\)

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle y' +y'\sin(x+y)= -\sin(x+y) }\)

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle y' (1+\sin(x+y))= -\sin(x+y) }\) \(\large\color{#000000 }{ \displaystyle y' =\frac{ -\sin(x+y) }{1+\sin(x+y)} }\)

OpenStudy (solomonzelman):

I can't read that code clearly, what is the point of tangency? )

OpenStudy (solomonzelman):

\((\pi/2~,~0)\) that?

OpenStudy (anonymous):

yeah

OpenStudy (solomonzelman):

So, plug that point \(x=\pi/2\) and \(y=0\) into the derivative, and see what you get.

OpenStudy (anonymous):

x=-1/2

OpenStudy (solomonzelman):

x ?

OpenStudy (solomonzelman):

you mean the result is -1/2?

OpenStudy (anonymous):

Oh I meant y'

OpenStudy (solomonzelman):

yeah :)

OpenStudy (solomonzelman):

-1/(1+1)=-1/2 (because sin(π/2)=1)

OpenStudy (solomonzelman):

So your slope is -1/2. And the point is \((\pi/2,0)\) you got it from here

OpenStudy (anonymous):

Ok, thank you very much!!

OpenStudy (solomonzelman):

If you would like to later confirm the answer, then, here is the graph of the curve, the point, and the tangent line. https://www.desmos.com/calculator/hdtir0corc

OpenStudy (anonymous):

Alright, thanks! :)

OpenStudy (solomonzelman):

Good Luck!